The perimeter of a rectangle is 700yds. What are the dimensions of the rectangle if the length is 60yds more then width.

Answer by checkley79(3341)

P=2L+2W
L=W+60
700=2(W+60)+2W
700=2W+120+2W
700-120=4W
580=4W
W=580/4
W=145 yds.
L=145+60
L=205 yds.
Proof:
700=2*205+2*145
700=410+290
700=700